# Level-Set Parameters: Novel Representation for 3D Shape Analysis

### v1

@@ Abstract
3D shape analysis has been largely focused on traditional 3D representations of point clouds and meshes, but the discrete nature of these data makes the analysis susceptible to variations in input resolutions. Recent development of neural fields brings in level-set parameters from signed distance functions as a novel, continuous, and numerical representation of 3D shapes, where the shape surfaces are defined as zero-level-sets of those functions. This motivates us to extend shape analysis from the traditional 3D data to these novel parameter data. Since the level-set parameters are not Euclidean like point clouds, we establish correlations across different shapes by formulating them as a pseudo-normal distribution, and learn the distribution prior from the respective dataset. To further explore the level-set parameters with shape transformations, we propose to condition a subset of these parameters on rotations and translations, and generate them with a hypernetwork. This simplifies the pose-related shape analysis compared to using traditional data. We demonstrate the promise of the novel representations through applications in shape classification (arbitrary poses), retrieval, and 6D object pose estimation. Code and data in this research are provided at this github link.
@@ 1 Introduction
3D surfaces are traditionally represented as point clouds or meshes on digital devices for visualization and geometry processing. This convention results in the current predominance of those data in 3D shape analysis, though the discrete nature of point clouds and polygon meshes can make the analysis approaches susceptible to variations in data resolutions [1, 2, 3, 4, 5, 6]. The recent advancements in neural fields enable manifold surfaces to be continuously represented by the zero-level-set of their signed distance functions (SDFs) [7, 8, 9]. Specifically, SDFs compute a scalar field by mapping each coordinate $\mathbf{x}\in\mathbb{R}^{3}$ to a scalar $v\in\mathbb{R}$ using a deep neural network, where the scalar $v$ indicates the signed distance of a point to its closest point on the surface boundary. Let $\boldsymbol{\theta}$ be the optimized parameters of the neural network. The shape surface can be numerically represented as $\boldsymbol{\theta}$, which we refer to as the level-set parameters. A typical example of 3D surface representation with level-set parameters could be $[n_{x},n_{y},n_{z},d]$ for a 3D plane of $\mathbf{n}^{\intercal}\mathbf{x}+d=0$, where $\mathbf{n}=[n_{x},n_{y},n_{z}]^{\intercal}$ and $\|\mathbf{n}\|=1$. The level-set parameters bring in novel 3D data for continuous shape analysis.
To initiate the novel shape analysis, we first have to construct the level-set parameters for each shape independently and with good accuracy in the dataset of interest. This differs from most methods in the SDF research domain, which are reconstruction-oriented and use a shared decoder together with various latent codes to improve the surface quality [7, 10, 11, 12, 13, 14]. However, the level-set parameters do not conform to Euclidean or metric geometry, unlike 3D point clouds. This presents a critical challenge for establishing good correlations in the parameter data of different shapes. Our solution is to formulate those parameters in high dimension as a pseudo-normal distribution with expectation $\boldsymbol{\mu}$ and identity covariance matrix $\mathbf{I}$, that is,
@@
We associate $\Delta\boldsymbol{\theta}$ with each individual shape, while the parameter $\boldsymbol{\mu}$ is shared by all shapes to (i) align different shapes in the parameter space, and (ii) initialize the SDF networks of individual shapes. Previous works train per-category SDF initializations using all training shapes [15] or a particular shape [16] in the category, which are either computation-intensive or dependent on the specific shape chosen. In contrast, we seek an initialization that generalizes to all categories of the dataset. Therefore, we take the trade-off of [15, 16] to use a few shapes from each category to learn $\boldsymbol{\mu}$. Other than learning an initialization, there are also methods employing identical random settings to initialize [17], which we empirically show yield insufficient correlations among shapes.
@@
Level-set parameters have rarely been considered as a data modality in shape analysis. Ramirez et al. [17] had to rely on the traditional data (i.e., point cloud and meshes) to extract shape semantics from level-set parameters using an encoder-decoder, due to their suboptimal SDF initializations. Erkoc et al. [16] utilized a small SDF network for tractable complexity in shape diffusion, whereas a small SDF network can undermine the representation quality of level-set parameters for complex shapes. More importantly, both methods are limited to shape analysis in the reference poses, ignoring the important shape transformations such as rotation and translation. In contrast, we explore the level-set parameters with transformations and extend the shape analysis to be pose-related.
@@
As is known, traditional data necessitate exhaustive data augmentations for shape analysis because transformations impact each element (i.e., point or vertex) in the data [1, 5]. Many endeavors have been made to address this issue using equivariant neural networks [18, 19, 20, 21, 22]. However, for shapes represented by level-set parameters, transformations usually affect a typical subset of the parameters (e.g., those in the first layer of SDF), potentially simplifying the subsequent analysis. We propose to condition those subset parameters on rotations and translations and generate them with a hypernetwork, which facilitates the analysis further.
@@
To acquire experimental data, we construct level-set parameters for shapes in the ShapeNet [23] and Manifold40 [6, 24] datasets. We demonstrate the potential of the proposed data through applications in shape classification of arbitrary poses, shape retrieval, and 6D object pose estimation. In pose estimation, we consider the problem of estimating shape poses from their partial point cloud observations, given that the level-set parameters are provided. This is similar to a partial-to-whole registration task [25]. The main contributions of this work are summarized as below:
@@
  • We open-source our code and level-set parameter data for potential interest in the community.
@@ 2.1 Shape Analysis with Traditional 3D Data
Semantic Analysis. The traditional 3D data, point clouds and polygon meshes, represent shape surfaces discretely in Euclidean space. Point clouds, as orderless collections of points, result in early-stage MLP-based neural networks [1, 26, 27] to introduce permutation-invariant operations for semantic learning from such data. Graph convolutional networks [4, 28, 29, 30] enable point clouds to be processed with convolutional operations which handle spatial hierarchies of data better than MLPs. Transformer-based networks [31, 32, 33] treat each point cloud as a sequence of 3D points, offering another approach. Point clouds can also be voxelized into regular grids and processed by 3D convolutional neural networks [24, 34, 35, 36]. Polygon meshes, which are point clouds with edge connections on the shape surface, provide more information about the shape geometry. Different approaches have been proposed to learn shape semantics from meshes [5, 6, 37, 38].
Most of the aforementioned methods require exhaustive data augmentations, such as rotations and translations, to maintain effectiveness with transformed data. Different equivariant neural networks are developed to incorporate rotation and translation equivariance to address this issue [18, 19, 20, 21, 22]. In comparison, we study shape semantics with the level-set parameters, which simplifies the problem.
Geometric Analysis. 6D object pose estimation is crucial for various applications such as robotics grasping [39], augmented reality [40], and autonomous driving [41]. It requires accurately determining the rotation and translation of an object relative to a reference frame. Despite the predominance of estimating object poses from RGB(D) data [42, 43, 44, 45], this task can alternatively be defined as a registration problem based on point cloud data. It involves registering the partial point cloud of an object (observation) to its full point cloud (reference) [25, 46]. We are interested in the potential of level-set parameters in the geometric analysis of 3D shapes and therefore represent the reference objects using level-set parameters instead of full point clouds. Leveraging the SDF reconstruction loss, we propose a registration method that requires no training data [47], correspondences [48, 49, 50, 51] or global shape features [52, 53]. We compare it with other optimization-based registration algorithms, including ICP [54], FGR [55], and TEASER(++) [56]. Go-ICP [57] is excluded here for complexity.
@@ 2.2 Neural Fields for 3D Reconstruction
Neural fields utilize coordinate-based neural networks to compute signed distance fields [7, 8] or occupancy fields [10] for 3D reconstructions. Many methods in this domain employ a shared decoder to reconstruct 3D shapes across an entire dataset or a specific category using different shape latent codes [7, 10, 11, 12, 58, 59]. The modelling function can be denoted as $f_{\boldsymbol{\theta}}:\mathbb{R}^{3}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$, with $\mathbf{z}\in\mathbb{R}^{m}$ being the shape latent code, which is learned by an encoder from various inputs [10], or randomly initialized and optimized as in DeepSDF [7]. Some works extend the concept of latent codes from per-shape to per-point for improved surface quality [13, 14]. In contrast, SIREN [8] applies an SDF network to instance-level surface reconstruction using periodic activations, which does not involve latent codes. It employs an unsupervised loss function, eliminating the need for ground-truth SDF values as in [7]. We note that neural radiance fields (NeRFs) [60, 61, 62, 63] reconstruct 3D surfaces with entangled neural parameters for surface geometry and photometry, where the latter is view-dependent. We therefore focus on surface geometry and utilize the level-set parameters from SDFs in our study. Our SDF network is adapted from the common 8-layer MLP utilized by [7, 58, 59].
@@ 2.3 Level-Set Parameters as 3D Data
Few works have studied the level-set parameters as an alternative data modality for 3D research, and each has utilized a different SDF network. Ramirez et al. [17] initialized the SDF network of SIREN [8] with identical random settings and trained the parameters independently for each shape, resulting in insufficient shape correlations in the parameter space. To extract shape semantics from the level-set parameters, theyo utilized an encoder-decoder architecture with additional supervision from traditional data. In contrast, our proposed method constructs the parameter data with improved shape correlations, enabling the learning of shape semantics using an encoder without relying on traditional data. We note that the periodic activation functions of SIREN cause undesired shape artifacts in empty spaces compared to ReLU activations [64]. Erkoc et al. [16] utilized a much smaller SDF network with ReLU activations for continuous shape generation. They initialized the network using overfitted parameters of a particular shape and trained the parameters of each shape within the same category independently. Yet, small SDF networks cannot represent complex shapes with good accuracy. Dupont et al. [65] aimed to explore the parameters of diverse neural fields as continuous data representations, but instead resorted to their modulation vectors [66, 67] for simplicity. They exploited the strategy of meta-learning [15, 68, 69] to construct those modulation vectors.
@@ 3 Dataset of Level-Set Parameters
Preliminaries. We adopt the well-established SDF architecture from previous research [7, 58], which comprises an 8-layer MLP with a skipping concatenation at the 4th layer. We utilize $256$ neurons for all interior layers other than the skipping layer which has $253$ neurons due to the input concatenation. Note that shape latent codes are not required. We employ smoothed ReLU as the activation function and train the SDF network $f(\mathbf{x};{\boldsymbol{\theta}})$ with the unsupervised reconstruction loss of SIREN [8], i.e.,
@@
$\mathcal{L}_{\text{dist}}^{p},\mathcal{L}_{\text{dist}}^{n}$ are the respective distances of positive and negative points to the surface. $\mathcal{L}_{\text{eik}}$ is the Eikonal loss [59] and $\mathcal{L}_{\text{norm}}^{p}$ imposes normal consistency. The constants $\lambda_{1\sim 4}$ balance different objectives.
@@ 3.1 Surface Transformation
Instead of generating the first layer parameters directly as $h(\mathbf{R},\mathbf{t};\boldsymbol{\eta})$, we introduce the latent matrices $\{\mathbf{Y}^{mn}\}$ to guarantee that our generated SDF parameters satisfy the geometric initializations recommended by SAL [58]. This is critical for the network convergence and good performance.
Generally, the hypernetwork calculates a compact matrix $\mathcal{B}{=}h(\mathbf{R},\mathbf{t};\boldsymbol{\eta})$ of size $256{\times}4{\times}I{\times}J$ according to $\mathbf{R},\mathbf{t}$, which contains the pose-dependent coefficient matrices $\mathbf{B}^{mn}$ associated with each latent matrix $\mathbf{Y}^{mn}$. Each pair of matrices $\mathbf{B}^{mn}$ and $\mathbf{Y}^{mn}$ is combined to compute a vector $\mathbf{z}^{mn}$, normalized into $\hat{\mathbf{z}}^{mn}$ to satisfy the standard normal distribution. In the last layer of $h_{\boldsymbol{\phi}}$, two branches of fully connected layers accept $\mathbf{z}^{mn}$ and $\hat{\mathbf{z}}^{mn}$, respectively, to initialize the biases and weights in $\{\theta_{1}^{mn}\}$ accordingly. See supplementary for the details. We note that the tanh activation in the final computation of $\mathcal{B}$ (see Fig. 1) helps to constrain all of its values within the range of $[-1,1]$.

### v2

@@ Abstract
3D shape analysis has been largely focused on traditional 3D representations of point clouds and meshes, but the discrete nature of these data makes the analysis susceptible to variations in input resolutions. Recent development of neural fields brings in level-set parameters from signed distance functions as a novel, continuous, and numerical representation of 3D shapes, where the shape surfaces are defined as zero-level-sets of those functions. This motivates us to extend shape analysis from the traditional 3D data to these novel parameter data. Since the level-set parameters are not Euclidean like point clouds, we establish correlations across different shapes by formulating them as a pseudo-normal distribution, and learn the distribution prior from the respective dataset. To further explore the level-set parameters with shape transformations, we propose to condition a subset of these parameters on rotations and translations, and generate them with a hypernetwork. This simplifies the pose-related shape analysis compared to using traditional data. We demonstrate the promise of the novel representations through applications in shape classification (arbitrary poses), retrieval, and 6D object pose estimation.
@@ 1 Introduction
3D surfaces are traditionally represented as point clouds or meshes on digital devices for visualization and geometry processing. This convention results in the current predominance of those data in 3D shape analysis, though the discrete nature of point clouds and polygon meshes can make the analysis approaches susceptible to variations in data resolutions Qi et al. (2017a); Wang et al. (2019a); Li et al. (2018); Lei et al. (2020); Hanocka et al. (2019); Hu et al. (2022). The recent advancements in neural fields enable manifold surfaces to be continuously represented by the zero-level-set of their signed distance functions (SDFs) Park et al. (2019a); Sitzmann et al. (2020a); Xie et al. (2022). Specifically, SDFs compute a scalar field by mapping each coordinate $\mathbf{x}\in\mathbb{R}^{3}$ to a scalar $v\in\mathbb{R}$ using a deep neural network, where the scalar $v$ indicates the signed distance of a point to its closest point on the surface boundary. Let $\boldsymbol{\theta}$ be the optimized parameters of the neural network. The shape surface can be numerically represented as $\boldsymbol{\theta}$, which we refer to as the level-set parameters. A typical example of 3D surface representation with level-set parameters could be $[n_{x},n_{y},n_{z},d]$ for a 3D plane of $\mathbf{n}^{\intercal}\mathbf{x}+d=0$, where $\mathbf{n}=[n_{x},n_{y},n_{z}]^{\intercal}$ and $\|\mathbf{n}\|=1$. The level-set parameters bring in novel 3D data for continuous shape analysis.
To initiate the novel shape analysis, we first have to construct the level-set parameters for each shape independently and with good accuracy in the dataset of interest. This differs from most methods in the SDF research domain, which are reconstruction-oriented and use a shared decoder together with various latent codes to improve the surface quality Park et al. (2019a); Mescheder et al. (2019); Erler et al. (2020); Chen and Zhang (2019); Peng et al. (2020); Chibane et al. (2020). However, the level-set parameters do not conform to Euclidean or metric geometry, unlike 3D point clouds. This presents a critical challenge for establishing good correlations in the parameter data of different shapes. Our solution is to formulate those parameters in high dimension as a pseudo-normal distribution with expectation $\boldsymbol{\mu}$ and identity covariance matrix $\mathbf{I}$, that is,
@@
We associate $\Delta\boldsymbol{\theta}$ with each individual shape, while the parameter $\boldsymbol{\mu}$ is shared by all shapes to (i) align different shapes in the parameter space, and (ii) initialize the SDF networks of individual shapes. Previous works train per-category SDF initializations using all training shapes Sitzmann et al. (2020b) or a particular shape Erkoç et al. (2023) in the category, which are either computation-intensive or dependent on the specific shape chosen. In contrast, we seek an initialization that generalizes to all categories of the dataset. Therefore, we take the trade-off of Sitzmann et al. (2020b); Erkoç et al. (2023) to use a few shapes from each category to learn $\boldsymbol{\mu}$. Other than learning an initialization, there are also methods employing identical random settings to initialize Luigi et al. (2023), which we empirically show yield insufficient correlations among shapes.
@@
Level-set parameters have rarely been considered as a data modality in shape analysis. Ramirez et al. Luigi et al. (2023) had to rely on the traditional data (i.e., point cloud and meshes) to extract shape semantics from level-set parameters using an encoder-decoder, due to their suboptimal SDF initializations. Erkoc et al. Erkoç et al. (2023) utilized a small SDF network for tractable complexity in shape diffusion, whereas a small SDF network can undermine the representation quality of level-set parameters for complex shapes. More importantly, both methods are limited to shape analysis in the reference poses, ignoring the important shape transformations such as rotation and translation. In contrast, we explore the level-set parameters with transformations and extend the shape analysis to be pose-related.
@@
As is known, traditional data necessitate exhaustive data augmentations for shape analysis because transformations impact each element (i.e., point or vertex) in the data Qi et al. (2017a); Hanocka et al. (2019). Many endeavors have been made to address this issue using equivariant neural networks Deng et al. (2021); Esteves et al. (2018); Cohen et al. (2018); Thomas et al. (2018); Poulenard and Guibas (2021). However, for shapes represented by level-set parameters, transformations usually affect a typical subset of the parameters (e.g., those in the first layer of SDF), potentially simplifying the subsequent analysis. We propose to condition those subset parameters on rotations and translations and generate them with a hypernetwork, which facilitates the analysis further.
@@
To acquire experimental data, we construct level-set parameters for shapes in the ShapeNet Chang et al. (2015) and Manifold40 Hu et al. (2022); Wu et al. (2015) datasets. We demonstrate the potential of the proposed data through applications in shape classification of arbitrary poses, shape retrieval, and 6D object pose estimation. In pose estimation, we consider the problem of estimating shape poses from their partial point cloud observations, given that the level-set parameters are provided. This is similar to a partial-to-whole registration task Dang et al. (2022). The main contributions of this work are summarized as below:
@@ 2.1 Shape Analysis with Traditional 3D Data
Semantic Analysis. The traditional 3D data, point clouds and polygon meshes, represent shape surfaces discretely in Euclidean space. Point clouds, as orderless collections of points, result in early-stage MLP-based neural networks Qi et al. (2017a); Klokov and Lempitsky (2017); Qi et al. (2017b) to introduce permutation-invariant operations for semantic learning from such data. Graph convolutional networks Lei et al. (2020); Wang et al. (2019b); Wu et al. (2019); Thomas et al. (2019) enable point clouds to be processed with convolutional operations which handle spatial hierarchies of data better than MLPs. Transformer-based networks Zhao et al. (2021); Guo et al. (2021); Wu et al. (2022) treat each point cloud as a sequence of 3D points, offering another approach. Point clouds can also be voxelized into regular grids and processed by 3D convolutional neural networks Wu et al. (2015); Maturana and Scherer (2015); Riegler et al. (2017); Graham et al. (2018). Polygon meshes, which are point clouds with edge connections on the shape surface, provide more information about the shape geometry. Different approaches have been proposed to learn shape semantics from meshes Hanocka et al. (2019); Hu et al. (2022); Smirnov and Solomon (2021); Lei et al. (2023).
Most of the aforementioned methods require exhaustive data augmentations, such as rotations and translations, to maintain effectiveness with transformed data. Different equivariant neural networks are developed to incorporate rotation and translation equivariance to address this issue Deng et al. (2021); Esteves et al. (2018); Cohen et al. (2018); Thomas et al. (2018); Poulenard and Guibas (2021). In comparison, we study shape semantics with the level-set parameters, which simplifies the problem.
Geometric Analysis. 6D object pose estimation is crucial for various applications such as robotics grasping Tremblay et al. (2018), augmented reality Marchand et al. (2015), and autonomous driving Geiger et al. (2012). It requires accurately determining the rotation and translation of an object relative to a reference frame. Despite the predominance of estimating object poses from RGB(D) data Wang et al. (2019c, d); Peng et al. (2019); Park et al. (2019b), this task can alternatively be defined as a registration problem based on point cloud data. It involves registering the partial point cloud of an object (observation) to its full point cloud (reference) Dang et al. (2022); Jiang et al. (2023). We are interested in the potential of level-set parameters in the geometric analysis of 3D shapes and therefore represent the reference objects using level-set parameters instead of full point clouds. Leveraging the SDF reconstruction loss, we propose a registration method that requires no training data Zeng et al. (2017), correspondences Choy et al. (2019); Wang and Solomon (2019); Huang et al. (2021); Ao et al. (2023) or global shape features Huang et al. (2020); Aoki et al. (2019). We compare it with other optimization-based registration algorithms, including ICP Besl and McKay (1992), FGR Zhou et al. (2016), and TEASER(++) Yang et al. (2020). Go-ICP Yang et al. (2015) is excluded here for complexity.
@@ 2.2 Neural Fields for 3D Reconstruction
Neural fields utilize coordinate-based neural networks to compute signed distance fields Park et al. (2019a); Sitzmann et al. (2020a) or occupancy fields Mescheder et al. (2019) for 3D reconstructions. Many methods in this domain employ a shared decoder to reconstruct 3D shapes across an entire dataset or a specific category using different shape latent codes Park et al. (2019a); Mescheder et al. (2019); Erler et al. (2020); Chen and Zhang (2019); Atzmon and Lipman (2020); Gropp et al. (2020). The modelling function can be denoted as $f_{\boldsymbol{\theta}}:\mathbb{R}^{3}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$, with $\mathbf{z}\in\mathbb{R}^{m}$ being the shape latent code, which is learned by an encoder from various inputs Mescheder et al. (2019), or randomly initialized and optimized as in DeepSDF Park et al. (2019a). Some works extend the concept of latent codes from per-shape to per-point for improved surface quality Peng et al. (2020); Chibane et al. (2020). In contrast, SIREN Sitzmann et al. (2020a) applies an SDF network to instance-level surface reconstruction using periodic activations, which does not involve latent codes. It employs an unsupervised loss function, eliminating the need for ground-truth SDF values as in Park et al. (2019a). We note that neural radiance fields (NeRFs) Martin-Brualla et al. (2021); Wang et al. (2021); Yu et al. (2022); Yariv et al. (2021) reconstruct 3D surfaces with entangled neural parameters for surface geometry and photometry, where the latter is view-dependent. We therefore focus on surface geometry and utilize the level-set parameters from SDFs in our study. Our SDF network is adapted from the common 8-layer MLP utilized by Park et al. (2019a); Atzmon and Lipman (2020); Gropp et al. (2020).
@@ 2.3 Level-Set Parameters as 3D Data
Few works have studied the level-set parameters as an alternative data modality for 3D research, and each has utilized a different SDF network. Ramirez et al. Luigi et al. (2023) initialized the SDF network of SIREN Sitzmann et al. (2020a) with identical random settings and trained the parameters independently for each shape, resulting in insufficient shape correlations in the parameter space. To extract shape semantics from the level-set parameters, theyo utilized an encoder-decoder architecture with additional supervision from traditional data. In contrast, our proposed method constructs the parameter data with improved shape correlations, enabling the learning of shape semantics using an encoder without relying on traditional data. We note that the periodic activation functions of SIREN cause undesired shape artifacts in empty spaces compared to ReLU activations Ben-Shabat et al. (2022). Erkoc et al. Erkoç et al. (2023) utilized a much smaller SDF network with ReLU activations for continuous shape generation. They initialized the network using overfitted parameters of a particular shape and trained the parameters of each shape within the same category independently. Yet, small SDF networks cannot represent complex shapes with good accuracy. Dupont et al. Dupont et al. (2022) aimed to explore the parameters of diverse neural fields as continuous data representations, but instead resorted to their modulation vectors Mehta et al. (2021); Chan et al. (2021) for simplicity. They exploited the strategy of meta-learning Sitzmann et al. (2020b); Finn et al. (2017); Tancik et al. (2021) to construct those modulation vectors.
@@ 3 Dataset of Level-Set Parameters
Preliminaries. We adopt the well-established SDF architecture from previous research Park et al. (2019a); Atzmon and Lipman (2020), which comprises an 8-layer MLP with a skipping concatenation at the 4th layer. We utilize 256256256 neurons for all interior layers other than the skipping layer which has 253253253 neurons due to the input concatenation. Note that shape latent codes are not required. We employ smoothed ReLU as the activation function and train the SDF network $f(\mathbf{x};{\boldsymbol{\theta}})$ with the unsupervised reconstruction loss of SIREN Sitzmann et al. (2020a), i.e.,
@@
$\mathcal{L}_{\text{dist}}^{p},\mathcal{L}_{\text{dist}}^{n}$ are the respective distances of positive and negative points to the surface. $\mathcal{L}_{\text{eik}}$ is the Eikonal loss Gropp et al. (2020) and $\mathcal{L}_{\text{norm}}^{p}$ imposes normal consistency. The constants $\lambda_{1\sim 4}$ balance different objectives.
@@ 3.1 Surface Transformation
Instead of generating the first layer parameters directly as $h(\mathbf{R},\mathbf{t};\boldsymbol{\eta})$, we introduce the latent matrices $\{\mathbf{Y}^{mn}\}$ to guarantee that our generated SDF parameters satisfy the geometric initializations recommended by SAL Atzmon and Lipman (2020). This is critical for the network convergence and good performance.
Generally, the hypernetwork calculates a compact matrix $\mathcal{B}{=}h(\mathbf{R},\mathbf{t};\boldsymbol{\eta})$ of size $256{\times}4{\times}I{\times}J$ according to $\mathbf{R},\mathbf{t}$, which contains the pose-dependent coefficient matrices $\mathbf{B}^{mn}$ associated with each latent matrix $\mathbf{Y}^{mn}$. Each pair of matrices $\mathbf{B}^{mn}$ and $\mathbf{Y}^{mn}$ is combined to compute a vector $\mathbf{z}^{mn}$, normalized into $\hat{\mathbf{z}}^{mn}$ to satisfy the standard normal distribution. In the last layer of $h_{\boldsymbol{\phi}}$, two branches of fully connected layers accept $\mathbf{z}^{mn}$ and $\hat{\mathbf{z}}^{mn}$, respectively, to initialize the biases and weights in $\{\theta_{1}^{mn}\}$ accordingly. See supplementary for the details. We note that the tanh activation in the final computation of $\mathcal{B}$ (see Fig. 1) helps to constrain all of its values within the range of [1,1][-1,1][ - 1 , 1 ].